OpenStudy (anonymous):

Find the derivative of y with respect to x, t, or θ, as appropriate y = ln (cos (ln θ)) the answer is -(tan (lnθ)/θ) i just dont know how to get that

OpenStudy (anonymous):

So this is just a multiple chain rule so it becomes \[\ln(v(u))\]\[v = \cos(u)\] \[u = \ln\] so you get \[(1/\cos(\ln \theta) )* (-\sin(\ln \theta) * (1/\theta)\] \[= -\tan(\ln \theta)/\theta \]

OpenStudy (anonymous):

thank you